### v2

# Multi-Relational Algebra and Its Applications to Data Insights

Abstract
A range of data insight analytical tasks involves analyzing a large set of tables of different schemas, possibly induced by various groupings, to find salient patterns. In particular, such analyses are about many-to-many transformations of tables, while the classic relational algebra is about one-to-one or many-to-one transformations. This paper presents Multi-Relational Algebra, which extends relational algebra for such transformations and their compositions. Multi-Relational Algebra introduces MultiRelation to model of a set of tables of different schemas. Importantly, while the information unit in Relational Algebra is a tuple, the information unit in Multi-Relational Algebra is a slice, which formally is a pair $(r,X)$ where $r$ is a (region) tuple, and $X$ is a (feature) table [^1] In this paper, “table” and “relation” are interchangeable.. Multi-Relational Algebra introduces three new fundamental algebraic operators, MultiSelect, MultiProject, and MultiJoin, which lift their counterparts Select, Project, and Join to transform MultiRelation to MultiRelation. Through various examples, we show that multi-relational algebra can effortlessly express many complex analytic problems, some of which are traditionally considered out of scope for relational analytics. We have implemented and deployed a service for multi-relational analytics. Due to a unified logical design, we are able to conduct systematic optimization for a variety of seemingly different tasks. Our service has garnered interest from over a hundred internal teams who have developed data-insight applications using it, and serves millions of operators daily.
1 Introduction
A range of data insight analytical tasks involves analyzing a large set of tables of different schemas, possibly induced by various groupings, to find salient patterns. A common starting point is a GROUP-BY GROUPING SETS query on a table, which prepares aggregated data for different groupings. An example is given in Figure 1. Note that query results are padded by a special value ALL, since the classic Relational Algebra requires one to output one table.
(a) A base table with schema [Device, Browser, Date, Revenue].
SELECT
Device, Browser, Date, SUM(Revenue)
FROM
base_table
GROUP BY GROUPING SETS
[Date],
[Device, Date], [Browser, Date],
[Device, Browser, Date]
(b) A group-by-grouping-sets query where we group by Date; group by Device, Date; group by Browser, Date; and finally group by Device, Browser, Date. For each group by we aggregate Revenue by SUM.
(c) GROUPING-SETS results.
Figure 1: A common starting point: A GROUP-BY GROUPING SETS query on a base table. The four group-by created results on four different schemas: [Date, Revenue], [Device, Date, Revenue], [Browser, Date, Revenue], and [Device, Browser, Date, Revenue]. We use a special ‘ALL’ values to pad these four tables into one.
A first observation of this paper is that in many of our use cases, users want to analyze the results of the GROUP-BY GROUPING SETS query in a way that respects the grouping structure. Specifically, note that the GROUP BY GROUPING SETS query uses the ALL-padding to glue tables of four different schemas: (1) [Date], (2) [Device, Date] (3) [Browser, Date], and (4) [Device, Browser, Date]. This padding alters the grouping structure, so for example, if one wants to recover the group of [Device, Date], one needs to use a predicate ‘Device != NULL AND Browser = NULL AND Date != NULL’. Therefore, to analyze different groupings separately, one needs to apply different predicates, which is cumbersome in view of composing analysis. From this, we argue that we should instead allow users to logically work with a set of multiple tables with different schemas. Specifically, we call the query results of GROUP-BY GROUPING SETS before the padding, as illustrated on the left side of Figure 2, a MultiRelation.
Figure 2: $\mathsf{multi\_relation}$. A MultiRelation is a data object which contains several tables of different schemas. We highlighted the table identified by the schema [Device, Date, SumRevenue] using a red box.
A second observation of this paper concerns the information unit of analysis. In Relational Algebra, the information unit is mostly a tuple, and in GROUP-BY this is extended to a set of tuples with one column. However, in many of our use cases, the information unit of analysis is a slice, which formally is captured by a pair $(r,X)$, where $r$ is a tuple, and $X$ is a table that can have multiple columns. We call $r$ region, and $X$ features. Further, a useful analysis often needs to analyze multiple slices at once. We call such an analysis on multiple slices a slice transformation. From this, we argue that one should design new algebraic operators, which analyze slices in a multi-relation using slice transformations. Further, these operators output a multi-relation, and thus enables composition of multiple such operators for more complex MultiRelation-to-MultiRelation transformations. Let us consider an analysis journey as an example:
Slice Outlier Analysis(Figure 3(a)): For each of the groupings (e.g., [Device, Date]), we want to partition by the non-Date columns, and detect timeseries outliers on the resulted [Date, SumRevenue] tables.
(a) Outlier analysis task on multi_relation.
MultiSelect(
multi_relation,
region_spaces=[
[Device], [Browser], [Device, Browser]
slice_transformations=[
TimeSeriesOutlierDetect(Date, SumRevenue)
predicates=[max_daily_revenue > 100,]
(b) A MultiSelect on $\mathsf{multi\_relation}$. Note that, different from the group-by grouping sets query where Date appears in every grouping set, the region spaces do not have Date, and the Date attribute appears as a feature for the slice transformation TimeSeriesOutlierDetect.
Figure 3: Analysis task and MultiSelect operator. A MultiSelect operator (Definition 7) analyzes slices in a MultiRelation using a slice transformation TimeSeriesOutlierDetect. For example, the region space [Device], combined with the feature set [Date, SumRevenue], matches the relation with schema [Device, Date, SumRevenue], which we then partition to generate slices and analyze using TimeSeriesOutlierDetect. Figure 4 gives more details of slice analysis with reference features.
In this case, the region tuple $r$ has schema [Device], the feature table $X$ has schema [Date, Revenue], and a timeseries outlier analysis needs to be applied to a feature table with two columns. Further, for any particular region say (Device=Pixel), it is useful to compare the timeseries with the timeseries at the top (i.e., the relation with schema [Date, Revenue] in multi_relation). Figure 4 illustrates such an analysis of two slices, which is named TimeSeriesOutlierDetect.
Figure 4: Slice processing. ‘RegionSpace + FeatureSet’ gives a schema from which we match a relation in multi_relation to work on. A (region_tuple, partitioned_table) pair is called a slice. Moreover, reference slice is obtained with region_space=[] and feature_set=$T$=[Date, SumRevenue] (Definition 9). TimeSeriesOutlierDetect analyzes both slices.
The first MultiSelect. A core operator for analyzing a multi-relation is MultiSelect (Definition 7), and we illustrate its application for this analysis task in Figure 3(b). For this multi-select, we combine each region space with the feature set [Date, SumRevenue] of the slice transformation to match a relation in multi_relation, and then partition by the region space to generate slices, and then use the slice transformation to perform the region-reference analysis (Figure 4).
MultiSelect is not just brute-force. The above may seem to suggest that MultiSelect is simply brute-force. However, there can be much faster evaluation, since evaluation of different region spaces can be correlated. As an example, note that there is a natural partial order on regions: for regions $\gamma_{1},\gamma_{2}$, we say $\gamma_{1}\prec\gamma_{2}$ if the first region is more fine-grained (e.g., (Device=Pixel, Browser=Chrome) $\prec$ (Device=Pixel)). Now, we observe that max_daily_revenue respects the region partial order. That is, if (Device=Pixel) has maximal daily revenue $\leq 100$, then (Device=Pixel, Browser=Chrome) must also, so we can skip any region smaller than (Device=Pixel), which can potentially save a lot of expensive outlier detection calls. Our discussion indicates that MultiSelect has a character of searching in a space of slices, which indicates a rich space of optimizations.
The second MultiSelect. Back to the analysis journey, for the output of the first MultiSelect, slice results are coalesced into another MultiRelation (Figure 5(a)), on which we can do further analysis. For example, a follow-up analysis may compute the total interestingness, by summing over dates, which corresponds to another MultiSelect (Figure 5(b)).
(a) multi_relation2. The output of the first MultiSelect specified in Figure 3(b). ‘interestingness’ column is an output of ‘TimeSeriesOutlierDetect’.
MultiSelect(
multi_relation2,
region_spaces=[
[Device], [Browser], [Device, Browser]
slice_transformations=[
SUM(interestingness)
(b) A second MultiSelect on multi_relation2. For respective relation in the multi-relation, we compute the sum of ‘interestingness’.
A summary. The aforementioned journey serves as a basic example of numerous use cases that exhibit a common pattern. To systematically address these diverse scenarios, we introduce a novel algebraic theory. This theory provides a foundational framework for answering a wide range of questions related to these use cases. Guided by this theoretical approach, we have developed a service capable of efficiently solving these analytical tasks. Specifically:
Multi-Relational Algebra. We describe (Section 2) Multi-Relational Algebra, which extends the Relational Algebra. Specifically, we define MultiRelation which models a set of slices, and define three new algebraic operators, MultiSelect, MultiProject, and MultiJoin, which lift their classic counterparts, Select, Project, Join, to transform MultiRelation to MultiRelation. Multi-Relational Algebra significantly broadens the analytical scope of relational analytics.
Applications of Multi-Relational Algebra. We have built a service for executing a pipe of multi-relational operators. Our service has garnered interest from over a hundred internal teams who have developed data-insight applications using it, and serves millions of operators daily (Figure 9). Section 4 reports some major applications.
2 Multi-Relational Algebra
In this section we describe the Multi-Relational Algebra. Table 1 gives an overview of the algebra with a side-by-side comparison with classic Relational Algebra.
Relational Algebra Multi-Relational Algebra
Relation: A set of tuples MultiRelation: A set of slices. A slice is defined as a pair $(r,X)$, where $r$ is a tuple, and $X$ is a table
Select: Transform and filter a set of tuples MultiSelect: Transform and filter a set of slices
Project: Restrict a relation to a set of attributes MultiProject: Restrict a multi-relation to a set of schemas
Join: Merge two relations into one, by merging tuple pairs $(t,t^{\prime})$ MultiJoin: Merge two multi-relations into one, by merging slice pairs $(s,s^{\prime})$, where $s=(r,X),s^{\prime}=(r^{\prime},X^{\prime})$
Table 1: An overview of Multi-Relational Algebra. This table gives a side-by-side comparison of Multi-Relational Algebra with classic Relational Algebra.
2.1 Data Model
We start with the definition of MultiRelation, which is a data object that packs a set of tables of different schemas. In this form, a multi-relation is not yet explicitly a set of slices. We discuss slice representation of a multi-relation later.
Definition 1 (MultiRelation).
Let $U$ be a universe of attributes. A MultiRelation $\Psi$ is a data object that consists of a set of relations of different relation schemas, where each relation schema is a subset of $U$.
With this definition, a usual relation can be recognized as a multi-relation where there is only a single schema.
Slice representation of a multi-relation. In relational algebra, the unit of information processing is a tuple. In multi-relational algebra, the information unit becomes a slice. That is, operators in the multi-relational algebra are about processing slices encoded in a multi-relation. We start with the definition of slice.
Definition 2 (Slice).
A slice is a pair $(r,X)$, where $r$ is a tuple called region, and $X$ is a table called features, such that $schema(r)\cap schema(X)=\emptyset$, where $schema(r)$ and $schema(X)$ are the list of attributes in $r$ and $X$, respectively.
We stress that a multi-relation as defined above does not explicitly encode slices. Rather, we need to represent a multi-relation as a set of slices. We explain this below, and begin with the definition of block.
Definition 3 (Multi-Relational Block).
A multi-relational block, or simply block, is a pair of schema $(\Gamma,\operatorname{{\mathcal{}F}})$ where $\Gamma\cap\operatorname{{\mathcal{}F}}=\emptyset$.
Definition 4 (Representing a block in a multi-relation).
Let $\Psi$ be a multi-relation, and $(\Gamma,\operatorname{{\mathcal{}F}})$ be a block. A slice representation of $\Psi$ with respect to block $(\Gamma,\operatorname{{\mathcal{}F}})$ is computed as follows:
  1. 1.
    Minimal Schema Matching. Let $\operatorname{{\mathcal{}T}}=\Gamma\cup\operatorname{{\mathcal{}F}}$. We find the relation $Z$ in $\Psi$, such that $\operatorname{{\mathcal{}T}}\subseteq schema(Z)$ (it contains the schema $\operatorname{{\mathcal{}T}}$), and is minimal in the sense of the size of the schema $|schema(Z)|$ (i.e., the number of attributes).
  2. 2.
    Projecting to $\operatorname{{\mathcal{}T}}$. We project $Z$ to $\operatorname{{\mathcal{}T}}$, denoted as $Z|_{\operatorname{{\mathcal{}T}}}$.
  3. 3.
    Partition to get slices. Finally, we partition $Z|_{\operatorname{{\mathcal{}T}}}$ over $\Gamma$, and get a set of slices $(r,X)$, where $schema(r)=\Gamma$ and $schema(X)=\operatorname{{\mathcal{}F}}$.
For (1), if there are multiple relations satisfy that condition, there are multiple ways to handle tie (for example, we can break the tie arbitrarily, or simply report an error). We leave details of handling ties later.
An important use of representation is to represent a multi-relation as different sets of slices. Below is an example
Example 1.
$\Psi$[$\Gamma$=[Device]][$\operatorname{{\mathcal{}F}}$=[Date, Revenue]] gives a representation of slices where region has schema [Device], and the table has schema [Date, Revenue]. The result is shown in Figure 6(a). Now Consider $\Psi$[$\Gamma$=[Device, Date]][$\operatorname{{\mathcal{}F}}$=[Revenue]]. This identifies slices where region has schema [Device, Date], and the features have schema [Revenue]. The result is shown in Figure 6(b).
(a) region_space=[Device],
feature_set=[Date, Revenue].
(b) region_space=[Device, Date],
feature_set=[Revenue].
Before we proceed to discuss operators, we define a semi-lattice structure on tuples which turns out to be important when processing regions.
Definition 5 (Region Lattice).
We note a natural partial order among regions: For two regions $\gamma_{1}$ and $\gamma_{2}$, $\gamma_{1}\prec\gamma_{2}$, if every attribute value in $\gamma_{2}$ is also in $\gamma_{1}$. In other words, $\gamma_{1}$ is more fine-grained, while $\gamma_{2}$ is more coarse-grained (e.g., (State=CA, City=MTV) $\prec$ (State=CA)).
This lattice structure has two main uses in our current applications: (1) It is used in MultiJoin for a type of join condition on regions, and (2) It gives a way to optimize MultiSelect by leveraging Apriori property of slice transformation output on this lattice.
2.2 Operators
In this section we describe three main operators in Multi-Relational Algebra: MultiSelect, MultiProject, MultiJoin, which lift their classic counters Select, Project, Join to transform MultiRelation to MultiRelation. At the high level, all these algebraic operators transform a multi-relation in the following basic pattern:
  1. 1.
    From the operator parameters, we obtain multiple slice representations of a multi-relation. In the case of a MultiJoin, we need to obtain slice representations for both input multi-relations.
  2. 2.
    We apply slice transformations to appropriate slice representations to obtain output slice representations. A slice transformation can take multiple slices as input. In most of our use case, we use slice transformations which takes two slices as input, which we call the 2-slice transformations.
  3. 3.
    All output slice representations are merged into an output multi-relation.
2.2.1 MultiSelect
In the relational algebra, Select is to apply transformations to tuples and filter. In Multi-Relational Algebra, this is generalized to apply transformations to slices and filter. We first define the simplest transformations applying to slices, the 1-slice transformations.
Definition 6 (1-Slice Transformation).
A 1-slice transformation $M$ takes one slice as input, and outputs another slice. Specifically, $M$ works as follows:
  1. 1.
[…]
  2. 2.
    At the execution time, $M$ receives a slice, $(r,X)$, where $X$ is required to have schema $\operatorname{{\mathcal{}F}}$ (for example, generated using the slice representation described in Definition 4). In the output slice $(r^{\prime},X^{\prime})$ from the transformation $M$, both $r^{\prime}$ and $X^{\prime}$ can also have arbitrary schemas.
In our applications, we often use slice transformation where $r=r^{\prime}$, we call such slice transformation region invariant. 1-slice transformations are somewhat limited, since they will only amount to “local analysis”. We use this definition first since it is easier to set up the concept. We will later generalize and consider 2-slice transformation (i.e., which takes two slices as input), which are most useful for our applications.
Definition 7 (MultiSelect).
Let $\Psi$ be a multi-relation, $\Gamma_{1},\dots,\Gamma_{n}$ be $n$ region space schemas, $M_{1},\dots,M_{k}$ be $k$ slice transformations, where $M_{i}$ has feature set $\operatorname{{\mathcal{}F}}_{i}$, and results_predicates be predicates for filtering results. A MultiSelect is defined as
MultiSelect(
$\Psi$,
region_spaces=[$\Gamma_{1}$, ..., $\Gamma_{n}$]
slice_transformations=[$M_{1}$, ..., $M_{k}$],
results_predicates,
and is processed as follows: For each $i\in[n],j\in[k]$,
  1. 1.
    For each $\Gamma_{i}$ and $\operatorname{{\mathcal{}F}}_{j}$ (of model $M_{j}$) we compute the slice representation $\operatorname{{\mathcal{}S}}^{i}_{j}$ of $\Psi$ w.r.t. $(\Gamma_{i},\operatorname{{\mathcal{}F}}_{j})$.
  2. 2.
    We apply $M_{j}$ to all slices in $\operatorname{{\mathcal{}S}}^{i}_{j}$, and filter the output slices using results_predicates.
Example 2.
Let $\Psi$ be a multi-relation as specified in Figure 2, LocalTimeSeriesOutlierDetect be 1-slice transformation which detects timeseries outlier with feature set [Date, SumRevenue] (which computes a p_val for every date in a date range), results_predicates be ‘p_val < .001’, and region spaces be [Device] and [Browser]. The following MultiSelect performs outlier detection for [Date, SumRevenue] timeseries slices for each specified region space:
MultiSelect(
$\Psi$,
region_spaces=[[Device], [Browser]],
slice_transformations=[
LocalTimeSeriesOutlierDetect(
time_attr=Date,
metric_attr=SumRevenue
predicates=[p_val < .001,]
Specifically:
  • for region space [Device], we match the relation with schema [Device, Date, SumRevenue] in $\Psi$, then we partition by Device, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and finally filter the output slice using ‘p_val < .001‘.
  • for region space [Browser], we match the relation with schema [Browser, Date, SumRevenue] in $\Psi$, then we partition by Browser, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and finally filter the output slice using ‘p_val < .001‘.
MultiSelect generalizes GROUP-BY GROUPING SETS and WINDOW functions. If we recognize a usual relation as a multi-relation with a single schema, then we observe that MultiSelect generalizes GROUP-BY GROUPING SETS. This is because GROUP-BY GROUPING SETS is about slice transformations that aggregate a single column. Therefore, MultiSelect also trivially generalize GROUP-BY and CUBE [GBLP96], which are special cases of GROUP-BY GROUPING SETS. MultiSelect also generalizes WINDOW functions, since we can use a slice transformation working on both the window column and metric column, and computes window-based aggregation (a more detailed explanation for this is deferred to A). Finally, we note that MultiSelect is trivially more general than all these operators since it can work on multi-relations.
2-slice transformations and MultiSelect with reference features. We first define 2-slice transformations, which is a slice transformation that takes two slices as input.
Definition 8 (2-Slice Transformation).
A 2-slice transformation $M$ is similar to a 1-slice transformation, except that its input has two slices, with potentially different schemas. Specifically, it works as follows:
  1. 1.
    The program of $M$ has two feature set parameters $\operatorname{{\mathcal{}P}}_{1}$ and $\operatorname{{\mathcal{}P}}_{2}$, which can be instantiated as two different feature schemas $\operatorname{{\mathcal{}F}}_{1}$ and $\operatorname{{\mathcal{}F}}_{2}$.
  2. 2.
    At the execution time, $M$ receives two slices as input, $(r_{1},X_{1})$ and $(r_{2},X_{2})$, where $X_{i}$ is required to have schema $\operatorname{{\mathcal{}F}}_{i}$. The transformation can perform any transformation and outputs outputs a slice $(r^{\prime},X^{\prime})$.
Given a multi-relation $\Psi$, we define its reference features as feature tables from the empty region
Definition 9 (Reference Features).
Given a multi-relation $\Psi$, and a feature set $\operatorname{{\mathcal{}F}}$, the reference feature data is the slice with feature set $\operatorname{{\mathcal{}F}}$ at the empty region, that is, $\Psi$(region=(), feature_set=$\operatorname{{\mathcal{}F}}$).
The following version of MultiSelect, which uses a 2-slice transformation so as to compare a local slice of a region with a global, reference slice, is very useful for our applications.
Definition 10 (Extending MultiSelect with reference features).
It is the same as MultiSelect except the following:
  1. 1.
    We use 2-slice transformations, and further each 2-slice transformation $M$ uses the same feature set for both parameters, that is $\operatorname{{\mathcal{}F}}_{1}=\operatorname{{\mathcal{}F}}_{2}=\operatorname{{\mathcal{}F}}$, where $\operatorname{{\mathcal{}F}}$ is a feature set.
  2. 2.
    Upon execution, it receives two slices (from (2), both feature tables has schema $\operatorname{{\mathcal{}F}}$): A slice at a particular region, and another slice at the empty region. The slice transformation analyzes both and computes an output slice.
Basically, 2-slice transformation with reference features allows one to analyze a region by comparing it with some reference, which significantly broadens the analytical problems one can solve. Example 3 shows a 2-slice transformation for outlier detection. We give a few examples in Section 3.
Example 3.
In Example 2 we described LocalTimeSeriesOutlierDetect, which analyzes a single slice from partitioning a matched relation. With reference features, we can extend this to TimeSeriesOutlierDetect, which is described earlier in Figure 4. Specifically, beside the region slice, TimeSeriesOutlierDetect also receives the reference slice, which is the “top-level” [Date, SumRevenue] timeseries. With this reference slice, TimeSeriesOutlierDetect can do many additional analysis, such as checking whether the region slice has too small of total revenue, and is therefore not worth analysis.
2.2.2 MultiProject
The Project operator in the classic relational algebra restricts a relation to a set of attributes. In the multi-relational algebra, MultiProject restricts a multi-relation to a set of schemas. Formally, MultiProject applies the block representation procedure 4 to multiple blocks:
Definition 11 (MultiProject).
Let $\Gamma_{1},\dots,\Gamma_{n}$ be region space schemas, and $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{k}$ be feature sets, then the MultiProject operator
MultiProject(
$\Psi$,
region_spaces=[$\Gamma_{1},\Gamma_{2},\dots,\Gamma_{n}$],
feature_sets=[$\operatorname{{\mathcal{}F}}_{1},\operatorname{{\mathcal{}F}}2,\dots,\operatorname{{\mathcal{}F}}_{m}$],
is the union of blocks for all $(\Gamma_{i},\operatorname{{\mathcal{}F}}_{j})$.
With MultiSelect and MultiProject we can define CRAWL operator:
Definition 12 (CRAWL).
CRAWL operator is a composition of a MultiSelect, followed by a MultiProject.
In view of SQL, CRAWL is the multi-relational generalization of “SELECT … WHERE …”, where the WHERE clause is a Select operator, and the SELECT clause is a Project operator.
2.2.3 MultiJoin
In the classic join, we have two tuples $t,t^{\prime}$, and we have a set of columns shared by both tuples as join keys, then we merge these two tuples into one $\widetilde{t}$, subject to some conditions on the join keys. In multi-relational algebra, the situation is similar, except that we merge two slices $(r,X)$ and $(r^{\prime},X^{\prime})$ into one $(\widetilde{r},\widetilde{X})$, where $r,r^{\prime}$ act as join keys. and the merge is captured by a 2-slice transformation.
Definition 13 (MultiJoin).
MultiJoin is an operator as follows:
MultiJoin(
$\Psi_{1}$,
$\Psi_{2}$,
region_spaces=[$\Gamma_{1},\dots,\Gamma_{n}$],
region_join_conditions,
slice_transformations=[$M_{1},\dots,M_{k}$],
Suppose we associate a region $r$ in $\Psi_{1}$ with another region $r^{\prime}$ in $\Psi_{2}$ that are joinable w.r.t. the specified conditions, For any 2-slice transformation $M_{i}$, it has two features sets $\operatorname{{\mathcal{}F}}_{1}$, $\operatorname{{\mathcal{}F}}_{2}$, from which we can identify a slice $S=(r,X)$ and $S^{\prime}=(r^{\prime},X^{\prime})$, and so we can compute a slice using $M(S,S^{\prime})$. We repeat this for all models, and merge all output slices into a multi-relation.
One application of MultiJoin is the Inner-Outer Join and recurrent crawl, which we describe in Section 3.4.
3 Multi-Relational Analytics
It should not be a surprise that multi-relational operators with proper slice transformations can solve a wide-variety of problems. For example, MultiSelect with some trivial slice transformations capture the following two previous work:
  • DIFF (Abuzaid et al. (2021), which generalizes several data explanation engines). [AKS+21] defined $\sf DIFF$ operator, which receives two data frames, and differentiate these two data to find interesting regions using some simple statistics. One can encode this as a MultiSelect by using a column ‘is_test’, where $\sf True$/$\sf False$ identifies the two tables to differentiate. Moreover, computing the differentiation statistics can be wrapped into a slice transformation, which loads ‘is_test’ column as a feature.
  • AutoSlicer (Liu et al. 2022). [LRC22] proposed AutoSlicer to help understand deeper model performance in complex data. Their “search space” is exactly the region space in CRAWL and the analysis on slices can be easily captured by a region analysis model.
In the rest of this section, we give a few more examples of multi-relational analytics.
3.1 Frequent Itemset Mining
Frequent Itemset Mining (FIM [FIM]) is a classic data mining problem. In this section, we show that it can be captured by a MultiSelect with a slice transformation that has access to reference slice, but otherwise trivial. To this end, recall that the input to FIM is a set of transactions, where each transaction is a set of items. Figure 7(a) gives an illustration. A pattern is simply a set of items, and the support of a pattern is the fraction of all transactions that contains a particular pattern (the fraction is taken over the total number of transactions). FIM is to find patterns that have large support.
To solve this problem using MultiSelect, we encode the transactions as a relation, ‘encoding_df’, as illustrated in Figure 7(b). Such encoding can be achieved using common libraries (e.g., TransactionEncoder in mlextend [Ras18]). We attach a new column, ‘Count’, so that each row (i.e. a transaction) has a count 1. With this encoding, we observe that a pattern becomes a tuple where all attributes take value 1. Therefore, by treating all item attributes as region attributes in the MultiSelect, we can thus crawl these regions for finding frequent patterns. The slice transformation we use only requests ‘Count’ feature with a SUM aggregation, and
  • for a pattern region, such as (Beer=$\boldsymbol{1}$, Diaper=$\boldsymbol{1}$), the region slice (SUM(Count) for the region) is the number of transactions that contain {Beer, Diaper}, and
  • the reference slice (i.e., empty region ()) is the total number of transactions.
Dividing these two thus computes the support for a pattern. Several remarks: (1) Note that FIM is not traditionally considered a relational analytics problem, yet this example shows that it can be effortlessly captured by Multi-Relational Algebra, a natural extension of Relational Algebra. In particular, this example shows the power of having a slice transformation analyzing multiple slices (in this case, a region slice and a reference slice). (2) There are many optimizations we can do to evaluate this MultiSelect efficiently for this particular slice transformation. For example, the classic Apriori algorithm [Apr] can be implemented as a breadth-first (over regions) evaluation of the MultiSelect. A detailed discussion of this is however beyond the scope of this paper, which we leave as future work.
Id Transaction
1 { Beer, Diaper}
2 { Beer, Diaper }
3 { Rum, Whisky}
Beer Diaper Rum Whisky Count
$\boldsymbol{1}$ $\boldsymbol{1}$ 0 0 $\boldsymbol{1}$
$\boldsymbol{1}$ $\boldsymbol{1}$ 0 0 $\boldsymbol{1}$
0 0 $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{1}$
Table 2: (a): A transaction data set. (b): ‘encoding_df’ that encodes a transaction set, so that the “items” in the transaction data set become attributes in the encoding_df. The first three rows are encoding of the transactions of the first three transactions in (a).
3.2 Region Attribution via MultiSelect with Aumann-Shapley Slice Transformation
A common data insight problem is the region attribution problem: Consider again the base table given in Table 1(a), suppose we aggregate Revenue to date 2024-05-12 and date 2024-05-13, and found there is a big change, then we want to slice by Browser and Device and find which region contributes to this change. Of course, this problem is trivial for if the metric is Revenue, which is summable. However, it becomes a non-trivial problem if the metric is not summable.
Formally, let $\mu$ be a metric. for a region $\gamma$, we use $\mu^{\gamma}$ to denote the value of metric at the region. For the empty region (i.e., $\gamma=()$), we use $\mu^{p}$ to denote the population metric value. Suppose now that we divide the data into test and reference segments, where $\mu$ takes value $\mu^{p}_{t}$ in the test segment, and takes value $\mu^{p}_{c}$ in the reference segment, so there is a population metric change $\Delta^{p}=\mu^{p}_{t}-\mu^{p}_{c}$. the region attribution problem is defined as follows.
Definition 14 (Region Attribution Problem).
A region attribution $\operatorname{RAS}(\cdot)$ is a function that assigns a score to each region, and satisfies two properties:
  1. 1.
    (Completeness) For the empty region, $\operatorname{RAS}(())=\Delta^{p}$.
  2. 2.
    (Additivity) For any two disjoint regions $\beta,\gamma$, $\operatorname{RAS}(\beta\cup\gamma)=\operatorname{RAS}(\beta)+\operatorname{RAS}(\gamma)$.
Additivity is a natural requirement: For example, we have different devices and their attributions should add up to $\Delta^{p}$.
It turns out that this problem can be solved by a MultiSelect with a 2-slice transformation, which we call the Aumann-Shapley slice transformation, since it is an application of the Aumann-Shapley method ([SS11, AS15]) to this problem. Specifically, this transformation works on a region slice and the reference slice, where the reference slice is the data at the empty region (as we defined in Definition 9), which is what we want to perform attribution.
To describe Aumann-Shapley slice transformation, we need the concept of a Region-Ambient Metric Model: A region-ambient metric metric model encodes how to combine metrics from a region $\gamma$, and metrics of the ambient of the region $\gamma^{c}$ (i.e., data not in $\gamma$), to recover the population metric value. The metrics needed in a metric model may be more than $\mu$ itself. We give two examples of region-ambient metric models:
  • A summable metric (e.g., Revenue). In this case, a region-ambient metric model is simply $F(w,w^{\prime})=w+w^{\prime}$. This metric model contains two metrics: $w$ and $w^{\prime}$, where $w$ is the metric value in a region, and $w^{\prime}$ is the metric value in the ambient of the region.
  • A density metric (e.g., CostPerClick). A density metric if $\mu=W/S$ where both $W$ and $S$ are summable. In this case, a natural metric model is to have both $W$ and $S$, and has four parameters: $F(w,w^{\prime},s,s^{\prime})=(w+w^{\prime})/(s+s^{\prime})$. Here $w,s$ are numerator and denominator in a region, and $w^{\prime},s^{\prime}$ are numerator and denominator of the ambient.
Let $F$ be a metric model with parameters $z_{1},z_{2},\dots,z_{n}$, then one can then apply the Aumann-Shapley method to it, with test/reference values of all metrics as two end points $P_{0},P_{1}$. Specifically, we can consider a straight line connecting these two points: $h(t)=(1-t)P_{0}+tP_{1},t\in[0,1]$. Then, for $G(t)=F(h(t))$, we have
[…]
Therefore, for the attribution to $z_{i}$, we can define it as
[…]
Then we can define the region attribution as the sum of $\operatorname{RAS}(\gamma;z_{i})$ for appropriate $z_{i}$’s. We show this for both summable and density metrics below.
Summable metric. For a summable metric with metric model $F(w,w^{\prime})=w+w^{\prime}$, $z_{1}=w$, and $z_{2}=w^{\prime}$. For region attribution, we only care about the attribution to $w$, and so we define $\operatorname{RAS}(\gamma)=\operatorname{RAS}(\gamma;w)$. In this case, $P_{0}=(\mu^{\gamma}_{c},\mu^{p}_{c}-\mu^{\gamma}_{c})$ and $P_{1}=(\mu^{\gamma}_{t},\mu^{p}_{t}-\mu^{\gamma}_{t})$, the integral evaluates to
$\operatorname{RAS}(\gamma)=\operatorname{RAS}(\gamma;w)=\mu^{\gamma}_{t}-\mu^{\gamma}_{c}$ (2)
which matches our intuition that region attribution is simply region metric shift!
Density metric. For a density metric with metric model $F(w,w^{\prime},s,s^{\prime})=(w+w^{\prime})/(s+s^{\prime})$, the region attribution requires combining the attribution to $w$ and attribution to $s$, and so $\operatorname{RAS}(\gamma)=\operatorname{RAS}(\gamma;w)+\operatorname{RAS}(\gamma;s)$, which gives:
$\begin{split}&\operatorname{RAS}(\gamma)\\ =&\operatorname{RAS}(\gamma;w)+\operatorname{RAS}(\gamma;s)\\ =&\frac{(w_{t}^{r}-w_{c}^{r})(s_{t}^{p}-s_{c}^{p})-(w_{t}^{p}-w_{c}^{p})(s_{t}^{r}-s_{c}^{r})}{(s_{t}^{p}-s_{c}^{p})^{2}}\Big{(}\ln(s_{t}^{p})-\ln(s_{c}^{p})\Big{)}+\frac{s_{t}^{r}-s_{c}^{r}}{s_{t}^{p}-s_{c}^{p}}\Big{(}\frac{w_{t}^{p}}{s_{t}^{p}}-\frac{w_{c}^{p}}{s_{c}^{p}}\Big{)}\end{split}$ (3)
which is much less intuitive compared to (2). Note that reference metrics play an essential role in the formula.
Following this modeling, one can prove that the above attribution satisfies both of the completeness and additivity requirements in Definition 14. However, for complex metric models, the closed forms of integration are harder to derive, and we use JAX [BFH+18] to compute the integration numerically.
3.3 MultiSelect with Cross-Correlation Slice Transformation
In the previous case, we can explain a metric change using the Aumann-Shapley method if the metric admits a differentiable metric model. However, in many cases, such a metric model is too much to hope for. For example, suppose we want to understanding how Cost changes can be “attributed” to the Budget changes, then there is no clear analytical formula between Cost and Budget. In this case, a natural consideration is to use rank correlation [Ran] to study the statistical correlation between Cost and Budget. However, one common issue is that at the aggregated level, the correlation can be very weak, which gives no insight. This is where MultiSelect is useful, which slices the data into smaller regions, and considers correlation at the region level, which may lead to much more meaningful insights.
In the rest of this section, we give an example of slice transformation which analyzes correlation between metrics in the case where there is a test/reference split, and we only care about the correlation across reference and test periods, and we want to ignore correlations within each period. Figure 7 motivates such a consideration. In this example, a change of budget is made at the end of the reference period, and one sees that as we level up the budget, the cost also levels up (so a positive correlation). However, note that within each period, Cost and Budget have perfect negative correlation. This combination of positive and negative correlation results in a weak correlation factor using the classic rank correlation formula, since the classic formula considers data in both periods in one batch.
Figure 7: An important scenario for Cross Rank Correlation. In either the reference or the test period, Cost and Budget has perfect negative correlation (Budget increases as Cost decreases and vice versa). However, cross the two periods, Budget and Cost have perfect positive correlation: Budget increase gives Cost increase.
To solve this problem, we instead only consider pairs of points, where one point is from the reference period, and the other is from the test period. Formally, for the test period, we have $N$ data points of (Budget, Cost) pairs $U=\{(B^{t}_{i},R^{t}_{i})\}_{i=1}^{N}$. Similarly, for the reference period, we have $V=\{B^{c}_{j},R^{c}_{j}\}_{j=1}^{M}$. The cross rank correlation formula is given in Equation 4.
$\text{CrossRankCorr}(U,V)=\frac{\sum_{i\in[N],j\in[M]}\text{sgn}\left(\Big{(}R^{t}_{i}-R^{c}_{j}\Big{)}\Big{(}B^{t}_{i}-B^{c}_{j}\Big{)}\right)}{NM}$ (4)
3.4 Inner-Outer Join
Consider the metric correlation described above. In many cases, the metrics are from different multi-relations (for example, they are collected from different channels). Therefore, we need to join these multi-relations in order to analyze correlations. Now consider two slices $s=(r,X)$ and $s^{\prime}=(r^{\prime},X^{\prime})$, it is natural that we put equi-condition on the regions (i.e. $r=r^{\prime}$). In particular, this is an inner join, since if a region only appears on one side, there is no need to compute correlations for the region. Now suppose we have two slices $(r,X)$ and $(r,X^{\prime})$, where $X,X^{\prime}$ are two timeseries we need to join. At this point, there can be missing dates in either $X$ or $X^{\prime}$, and so we need a (full) outer join to join $X$ and $X^{\prime}$. We call such a case Inner-Outer Join. Moreover, since MultiJoin allows any transformation on $X,X^{\prime}$, we can use a composition of an outer join and a correlation detection on the resulting relation, expressed as follows using pipe syntax [Pyt]:
OuterJoin(key_attrs=[Date])
| CorrelationDetect(
entity_attrs=[Date],
metric1=’Revenue’,
metric2=’Budget’
The entire analysis can thus be expressed as a single MultiJoin
MultiJoin(
$\Psi_{1}$,
$\Psi_{2}$,
region_spaces=[$\Gamma_{1},\dots,\Gamma_{n}$],
region_join_conditions,
slice_transformations=[
OuterJoin(key_attrs=[Date])
| CorrelationDetect(
entity_attrs=[Date],
metric1=’Revenue’,
metric2=’Budget’
3.5 Recurrent CRAWL
We consider a case of composing CRAWL and MultiJoin for continual anomaly monitoring. Consider our earlier example of daily anomaly detection again, where we output a multi-relation with filtered daily p-values (i.e., p-value < .001). Let this multi-relation be $\Psi^{\prime}$. Now, on a new day, suppose we have another multi-relation $\Psi$ of daily-revenue, but we want to join with $\Psi^{\prime}$ and use the historical p-values as additional input for the analysis. In this case, we would then want to join $\Psi$ and $\Psi^{\prime}$ (with a left-outer join) to augment $\Psi$ with the p-value data in $\Psi^{\prime}$.
Transient Anomalies. An application of historical p-values is the use case of continual monitoring where there are transient anomalies. A transient anomaly is a data point that deviates significantly from norm for a very short period of time, and disappears quickly. Transient anomalies happen quite often in our data, and usually lead to false alarms.
Recurrent CRAWL. To tackle the transient anomalies, we use a multi-relational pipe called Recurrent CRAWL. Basically, at a new time point $t^{*}$, we have the current daily revenue multi-relation $\Psi$, and the multi-relation of historical p-value results $\Psi^{\prime}$. We first do a join as described earlier to merge them into $\widetilde{\Psi}$, and then perform a MultiSelect on $\widetilde{\Psi}$ with a slice transformation that examine both the revenue timeseries and the p-value timeseries. The specialty of this slice transformation is that it removes transient anomalies by replacing transient anomalous data values (as indicated by the historical p-values) with estimated normal values. Finally, it performs the usual outlier detection on the revised timeseries. Figure 8 visualizes the work of recurrent crawl.
Figure 8: Unrolling a recurrent CRAWL pipe. For each time point, we first join two the revenue multi-relation and the historical p-value multi-relation, then we apply a CRAWL on the resulting joined multi-relation. The historical p-values affects the new analysis at a time point (and thus the output p-value), and this effect carries on recursively.
4 Applications
We have built a service for executing a pipe of multi-relational algebra operators. Our service has attracted more than a hundred internal teams to build data insights applications with it. Figure 9 plots the daily usage of the service since Sept, 2023.
Figure 9: Daily number of operators served by our multi-relational analysis service. The x-axis is the time (date), and the y-axis is the number of operators we served for a particular day. The peak is about 12 million operators a day.
We briefly summarize three representative categories of applications:
  1. 1.
    System monitoring. Many of our clients have extremely complex systems, and monitoring the health of these systems is critical. These monitoring oftentimes require fine-grained understanding of the system, such as for a specific version of a system in a specific Geo, one has gradual increase of average latency.
  2. 2.
    Business intelligence (BI). Several of our major BI platforms need to generate data insights to help users better understand what is going on and identify new business opportunities. The requirements are often complex and ad-hoc, such as what campaigns and adgroups are affecting the top level Revenue change and how they are correlated with budget allocation.
  3. 3.
    Experimentation analysis. Several experimentation platforms have been built to slice the data to do A/B testing to understand the fine-grained impact of an experiment.
The multi-relational algebra pipes service provides high-level abstractions that benefit all these applications:
Our approach drastically simplifies system construction. For instance, in experimentation analysis platforms, teams often develop custom slicing-analysis kernels, each with tens of thousands of lines of code, typically built on Beam for its flexibility over SQL. Our multi-relational algebra pipes streamline these kernels into a concise pipeline of operators, enhancing interoperability with SQL due to their shared relational data foundation. This results in significant simplification of both experimentation analysis and integration with relational databases.
Enabling Advanced Capabilities. The composability of multi-relational operators has unlocked new and powerful functionalities. For example, in system monitoring, techniques like “recurrent crawl,” 3.5 built on these operators, have achieved superior precision and recall in anomaly detection compared to traditional statistical methods. This demonstrates the ability of multi-relational operators to express complex analysis patterns that would be difficult to achieve otherwise.
Enhanced Capabilities, Reduced Costs. The flexibility of slice transformations enables seamless integration of the Python Data Science stack with multi-relational processing. This allows for advanced analysis, such as quickly identifying revenue anomalies and their correlation with campaign budget allocation, leading to significant revenue gains in some cases. Additionally, migrating ad-hoc experimentation analysis systems to multi-relational pipes not only expands their capabilities to analyze finer-grained regions but also surprisingly results in improved latency, scalability, and reduced resource costs. This efficiency stems from the unified logical design of multi-relational algebra, enabling systematic optimizations across diverse analytical tasks.
5 Conclusion
From a theoretical perspective, the Join operator in Relational Algebra (RA) is unique in its ability to integrate multiple input tables into a single output table. We argue that this traditional approach is overly restrictive for the complex analytical tasks prevalent in modern data analysis. To address this limitation, we introduce Multi-Relational Algebra, an extension of RA that incorporates a novel data model called MultiRelation. This new model, along with accompanying operators, enables many-tables to many-tables transformations. Mutl-Relational Algebra significantly broadens the scope of algebraic theory for relational analytics, and has found various applications in data insights.
Appendix A Emulating WINDOW functions
A typical WINDOW function query is as follows:
SELECT
Name, Age, Department, Salary,
AVERAGE(Salary)
OVER( PARTITION BY Department ORDER BY Age)
AS AvgSalary
FROM
employee
For this query, we first partition the table by Department, so each partition is a table on schema [Name, Age, Salary]. Then for each partition, we sort by Age, and compute a moving average of Salary for each row. We can emulate this using MultiSelect as follows:
MultiSelect(
employee,
region_spaces=[[Department]],
slice_functions=[
MOVING_AVERAGE(
entity_attrs=[Name, Age], /* attributes for forming a row */
metric_attr=Salary,
order_by_attr=Age,
Note that we recognize the relation ‘employee’ as a multi-relation with a single schema.
Appendix B Extended operators
The operators described in the last section are fundamental, and already enabled a lot of otherwise-cumbersome analysis if one stays with RA. However, similar to that RA is extended with GroupBy, we consider two group-by extensions.
B.1 MultiGroupBy
In the RA, GroupBy partitions a relation into groups, and then uses an aggregation function to condense each group into a tuple, and finally we coalesce these tuples back into one relation. For BiRA, the situation is similar, we partition a bilevel relation into groups, where each group is a bilevel relation. Then we transform each group into another bilevel relation. Finally, we coalesce these bilevel relations into one by flattening the group key into the regions of each bilevel relation, and take union of all region-feature pairs.
Definition 15 (MultiGroupBy).
BiGroupBy is an operator as follows:
BiGroupBy(
br,
by_attributes,
bilevel_transform,
This works as follows:
  1. 1.
    We partition br using by_attributes.
  2. 2.
    For each group, we apply bilevel_transform to transform each bilevel relation to another.
  3. 3.
    We flatten the group keys into the regions of t bilevel relation, and take union of all region-feature pairs.
B.2 CoMultiGroupBy
Consider two bilevel relations br1 and br2, and we groupby for each on Country. Now, we want the bilevel relation in br1 under Country=US to co-occur with the bilevel relation in br2 under Country=US. Such a need of co-occurrence is not new: this also appears in Apache Beam CoGroupByKey [CoG]. However, unlike Beam which only needs to work with a data model of general dictionaries, here we must eventually output one bilevel relation. Suppose we use br1{Country=US} to denote the group Country=US after groupby, then co-groupby will put this together with br2{Country=US}, so we have a pair of bilevel relations (br1{Country=US}, br2{Country=US}). Then, we can use a BiJoin to condense this pair of bilevel relations back into one. After this, we are then in the situation of BiGroupBy, which we know how to handle.
Definition 16 (CoMultiGroupBy).
CoBiGroupBy is an operator as follows:
CoBiGroupBy(
br1,
br2,
by_attributes,
bijoin,
This works as follows:
  1. 1.
    We partition br1 and br2 by by_attributes to form groups, where each group is a bilevel relation.
  2. 2.
    For each group key K, we have a bilevel relation br1{K}, and another bilevel relation br2{K}.
  3. 3.
    We join these two bilevel relations into one, via bijoin(br1{K}, br2{K}).
After the above steps, we are in the same situation as BiGroupBy, which we handle accordingly.

### v3

@@
 A range of data insight analytical tasks involves analyzing a large set of tables of different schemas, possibly induced by various groupings, to find salient patterns.
-In particular, such analyses are about many-to-many transformations of tables, while the classic relational algebra is about one-to-one or many-to-one transformations.
-This paper presents Multi-Relational Algebra, which extends relational algebra for such transformations and their compositions.
-Multi-Relational Algebra introduces MultiRelation to model of a set of tables of different schemas.
-Importantly, while the information unit in Relational Algebra is a tuple, the information unit in Multi-Relational Algebra is a slice, which formally is a pair $(r,X)$ where $r$ is a (region) tuple, and $X$ is a (feature) table [^1] In this paper, “table” and “relation” are interchangeable..
-Multi-Relational Algebra introduces three new fundamental algebraic operators, MultiSelect, MultiProject, and MultiJoin, which lift their counterparts Select, Project, and Join to transform MultiRelation to MultiRelation.
-Through various examples, we show that multi-relational algebra can effortlessly express many complex analytic problems, some of which are traditionally considered out of scope for relational analytics.
+This paper presents Multi-Relational Algebra, an extension of the classic Relational Algebra, to facilitate such transformations and their compositions.
+Multi-Relational Algebra has two main characteristics: (1) Information Unit.
+The information unit is a slice $(r,X)$, where $r$ is a (region) tuple, and $X$ is a (feature) table [^1] In this paper, “table” and “relation” are interchangeable..
+Specifically, a slice can encompass multiple columns, which surpasses the information unit of “a single tuple” or “a group of tuples of one column” in the classic relational algebra, (2) Schema Flexibility.
+Slices can have varying schemas, not constrained to a single schema.
+This flexibility further expands the expressive power of the algebra.
+Through various examples, we show that multi-relational algebra can effortlessly express many complex analytic problems, some of which are beyond the scope of traditional relational analytics.
 We have implemented and deployed a service for multi-relational analytics.
 Due to a unified logical design, we are able to conduct systematic optimization for a variety of seemingly different tasks.
-Our service has garnered interest from over a hundred internal teams who have developed data-insight applications using it, and serves millions of operators daily.
+Our service has garnered interest from numerous internal teams who have developed data-insight applications using it, and serves millions of operators daily.
 1 Introduction
@@
 Multi-Relational Algebra.
-We describe (Section 2) Multi-Relational Algebra, which extends the Relational Algebra.
-Specifically, we define MultiRelation which models a set of slices, and define three new algebraic operators, MultiSelect, MultiProject, and MultiJoin, which lift their classic counterparts, Select, Project, Join, to transform MultiRelation to MultiRelation.
+Section 2 describes Multi-Relational Algebra, which extends Relational Algebra.
+Specifically, this algebra introduces MultiRelation for modeling a set of slices, and three new operators, MultiSelect, MultiProject, and MultiJoin, which lift their classic counterparts, Select, Project, Join, to transform MultiRelation to MultiRelation.
 Multi-Relational Algebra significantly broadens the analytical scope of relational analytics.
-Applications of Multi-Relational Algebra.
-We have built a service for executing a pipe of multi-relational operators.
-Our service has garnered interest from over a hundred internal teams who have developed data-insight applications using it, and serves millions of operators daily (Figure 9).
-Section 4 reports some major applications.
+For example, it captures two prior work, DIFF [ABD+18] and AutoSlicer [LRC22], both previously considered “beyond relational algebra.”
+Multi-Relational Analytics.
+We have built a service for constructing and evaluating multi-relational analytics.
+Our service has garnered interest from numerous internal teams who have developed applications using it, and serves millions of operators daily (Figure 9).
+Section 4 gives an overview of the service, and reports main application categories.
 2 Multi-Relational Algebra
@@
 Definition 3 (Multi-Relational Block).
-A multi-relational block, or simply block, is a pair of schema $(\Gamma,\operatorname{{\mathcal{}F}})$ where $\Gamma\cap\operatorname{{\mathcal{}F}}=\emptyset$.
+A multi-relational block, or simply block, is a pair of schemas $(\Gamma,\operatorname{{\mathcal{}F}})$, where $\Gamma\cap\operatorname{{\mathcal{}F}}=\emptyset$.
 Definition 4 (Representing a block in a multi-relation).
@@
 3 Multi-Relational Analytics
-It should not be a surprise that multi-relational operators with proper slice transformations can solve a wide-variety of problems.
-For example, MultiSelect with some trivial slice transformations capture the following two previous work:
+Multi-relational operators with slice transformations capture a wide-variety of problems, some of which are beyond the scope of traditional relational analytics.
+This power stems from two aforementioned characteristics of this algebra: (1) Information Unit.
+The information unit is a slice $(r,X)$, which can encompass multiple columns.
+This surpasses the information unit of “a single tuple” or “a group of tuples of one column” in the classic relational algebra, (2) Schema Flexibility.
+Slices can have varying schemas, not constrained to a single schema.
+This flexibility further expands the expressive power of the algebra.
+As initial illustrations, MultiSelect combined with simple slice transformations can express the functionality of two prior works, DIFF [ABD+18] and AutoSlicer [LRC22], both previously considered “beyond relational algebra.” DIFF, introduced as a novel relational operator, encompasses the capabilities of various ad-hoc data explanation engines.
+AutoSlicer, implemented using Apache Beam, serves in machine learning model validation.
   • DIFF (Abuzaid et al.
@@
 [LRC22] proposed AutoSlicer to help understand deeper model performance in complex data.
-Their “search space” is exactly the region space in CRAWL and the analysis on slices can be easily captured by a region analysis model.
+Their “search space” is exactly the region space in CRAWL and the analysis on slices can be easily captured by a 1-slice transformation (Definition 6).
 In the rest of this section, we give a few more examples of multi-relational analytics.
@@
 Dividing these two thus computes the support for a pattern.
-Several remarks: (1) Note that FIM is not traditionally considered a relational analytics problem, yet this example shows that it can be effortlessly captured by Multi-Relational Algebra, a natural extension of Relational Algebra.
+Remark 1.
+Several remarks:
+  1. 1.
+    Note that FIM is not traditionally considered a relational analytics problem, yet this example shows that it can be effortlessly captured by Multi-Relational Algebra, a natural extension of Relational Algebra.
 In particular, this example shows the power of having a slice transformation analyzing multiple slices (in this case, a region slice and a reference slice).
-(2) There are many optimizations we can do to evaluate this MultiSelect efficiently for this particular slice transformation.
+  2. 2.
+    There are many optimizations we can do to evaluate this MultiSelect efficiently for this particular slice transformation.
 For example, the classic Apriori algorithm [Apr] can be implemented as a breadth-first (over regions) evaluation of the MultiSelect.
-A detailed discussion of this is however beyond the scope of this paper, which we leave as future work.
+We briefly discuss this optimization in Section 4.2.
+  3. 3.
+    Astute readers may argue that our approach requires encoding transactions in a very wide table, with most of the table 00 (the number of columns is the number of distinct items), and it is wasteful to do this encoding.
+There is an elegant way to avoid such waste by exploiting the schema of a multi-relation, which we discuss in Section B.
 Id Transaction
@@
 The historical p-values affects the new analysis at a time point (and thus the output p-value), and this effect carries on recursively.
-4 Applications
+4 Multi-Relational Pipe and Applications
 We have built a service for executing a pipe of multi-relational algebra operators.
 Our service has attracted more than a hundred internal teams to build data insights applications with it.
-Figure 9 plots the daily usage of the service since Sept, 2023.
+Figure 9 plots the daily usage of the service since Sept, 2023, which has a daily peak of serving about 12 million operators.
+In this section we give a brief overview of our pipe service and its applications.
+Specifically, Section 4.1 presents an end-to-end example of a multi-relational pipe.
+Then, Section 4.2 briefly describes main optimizations for evaluating MultiSelect.
+Finally, we discuss main application categories in Section 4.3.
 Figure 9: Daily number of operators served by our multi-relational analysis service.
@@
 The peak is about 12 million operators a day.
+4.1 An example multi-relational pipe
+Consider a hypothetical scenario where we deploy a service on different devices and browsers, and we collect aggregated Cost and Clicks for each device-browser combinations.
+Figure 10 gives a hypothetical data.
+Suppose that an analyst wants to analyze the change of CostPerClick (Cpc) by comparing two set of data points: the data points where is_test=True, and the data point where is_test=False.
+The aggregated Cpc (SUM(Cost)/SUM(Clicks)) changes significantly: For is_test=True, the overall Cpc is about 15400.15, and for is_test=False, the overall Cpc is 56.25.
+Figure 10: pipe_df.
+A base table for finding insights.
+For is_test=True, CostPerClick = (100 + 100 + 2 + 200,000)/(1 + 1 + 1 + 10) $\approx$ 15400.15.
+For is_test=False, CostPerClick = (100 + 100 + 300 + 400)/(4 + 4 + 4 + 4) = 56.25.
+We want to find regions in Device and Browser that contribute to this change.
+The first question.
+As a first question, the analyst wants to know which regions in Device and Browser are the main contributors.
+To answer this question, we first filter rows where the metrics are NULL, and then do a CRAWL on the regions in Device and Browser (which contains 3 region spaces, [Device], [Browser], [Device, Browser]) and use density Aumann-Shapley attribution (described in Section 3.2) to analyze each region.
+This is expressed as a pipe in Listing 1.
+The ‘density_analysis’ computes the metric aumann_shapley_attribution, which we filter by ‘> .5’.
+The results of the first pipe are shown in Figure 11.
+We have found the two rows, row-2 and row-3, as the main contributors.
+This is encoded in two regions, (Device=Pixel) which corresponds to both row-2 and row-3, and (Device=Pixel, Browser=Edge), which corresponds to row-3.
+density_analysis = DensityDifferentiation(
+mass_column=’Cost’,
+volume_column=’Clicks’,
+is_test_column=’is_test’,
+results = (
+pipe_df
+| SqlTransform(’SELECT * FROM pipe WHERE Cost IS NOT NULL’)
+| CreateMultiTable(all_region_columns=[’Browser’, ’Device’, ’is_test’],)
+| Crawl(
+region_spaces=ALL_SPACES([’Browser’, ’Device’]),
+slice_transformations=[density_analysis],
+predicates=[aumann_shapley_attribution > 0.5,])
+Listing 1: A first pipe for finding important regions that contribute to a change. The slice transformation uses the density aumann-shapley attribution described in Section 3.2.
+Figure 11: Resulting of executing pipe 1.
+As expected (from eyeballing the data since it is small enough), we find row-2 and row-3 in table 10 as the main contributors, where row-2 drags the Cpc to be smaller, while row-3 drags the Cpc to be bigger.
+The region (Device=Pixel,) contains both row-2 and row-3, which have the biggest influence, the second region (Device=Pixel, Browser=Edge) contains only row-3, which is the biggest contributor to make the Cpc bigger for is_test=True.
+The second question.
+Note that Device=Pixel appears in both of the result regions, while Browser=Edge appears in only one.
+It is a common interest of our users to find frequent attribute patterns.
+This can be achieved using a second pipe, by crawling on the flattened results of the first analysis, using a frequent itemset analysis (as described in Section 3.1).
+This is expressed in Listing 2.
+The results of this pipe are shown in Figure 12: As expected, (Device=Pixel) is found as the most frequent pattern (support=1.0; ‘support’ of a pattern is defined in the first paragraph of Section 3.1).
+fim_analysis = FrequentItemsetAnalysis(
+count_column=’normalized_aumann_shapley_attribution’,
+results2 = (
+results
+| FlattenMultiRelation()
+| CreateMultiRelation(all_region_columns=[’Browser’, ’Device’])
+| Crawl(
+region_spaces=ALL_SPACES([’Browser’, ’Device’]),
+slice_transformations=[fim_analysis],
+Listing 2: A second pipe for finding frequent attribute patterns in the results of pipe 1. The slice transformation uses the frequent itemset slice transformation described in Section 3.1.
+Figure 12: Resulting of executing pipe 2. As expected, we find the Device=Pixel to be a frequent pattern with support 1, since it appears in all the regions in Figure 11.
+4.2 Optimizing MultiSelect with Apriori Property
+As we mentioned earlier, MultiSelect is not just brute-force, since evaluation of the slice transformations can be correlated between region spaces.
+Further, a unified logical modeling of MutltiSelect allows us to optimize for a variety of seemingly different analytical tasks.
+Some of these optimizations are beyond the scope of classical relational algebra, which is designed for processing set of tuples rather than set of slices.
+In this section we give a main example: Optimizing MultiSelect with Apriori property.
+To start with, recall the partial order of regions defined in Definition 5.
+Definition 15 defines Apriori property of slice transformation results w.r.t.
+this partial order:
+Definition 15 (Apriori property of slice transformation results).
+Let $\sigma$ be a metric computed by a slice transformation.
+The metric is said to satisfy the apriori property if for any two regions $\gamma_{1}\prec\gamma_{2}$, $\sigma[\gamma_{1}]\prec\sigma[\gamma_{2}]$ [^2] We use $\prec$ to denote a partial order for metrics, which may not be the linear order $<$ on real numbers., where $\sigma[\gamma]$ denotes the value of metric $\sigma$ on region $\gamma$.
+Apriori property and early stopping.
+If a slice transformation computes a metric that satisfies the Apriori property, then filtering on the metric allows us to early stop.
+This can be best illustrated on the CRAWL for frequent itemset mining (both Section 3.1 and pipe 2).
+For frequent itemset mining, the metric ‘support’ satisfies the Apriori property.
+For example, for the standard frequent itemset mining described in Section 3.1, if we filter by ‘support > .1’, then if the region (Beer=1,) already has support <= .1, then we can stop there without considering any region (pattern) that is more fine-grained than (Beer=1,).
+Degree-first evaluation and Apriori Algorithm for Frequent Itemset Mining.
+One can choose to leverage the Apriori property for early stopping in various ways.
+For example, one can do a degree-first333 the number of attributes in a region space is called the degree of the region space.
+evaluation of region spaces: We first consider evaluating all region spaces of degree 1 ([Beer], [Diaper], etc.), and then with the results of degree-1 regions, we filter by ‘support’, and only on those regions we consider region spaces of degree 2 ([Beer, Diaper], [Beer, Rum], etc.).
+This corresponds to a Breadth-First Search of regions.
+In fact, astute readers will recognize that this degree-first evaluation recovers the well known Apriori algorithm for Frequent Itemset Mining [Apr].
+This shows that our systematic optimization from considering Apriori property captures such an algorithm as a very special case.
+And the same optimization applies to many other tasks, as long as we filter by a metric with the Apriori property and do degree-first evaluation w.r.t.
+region spaces.
+Beyond degree-first evaluation.
+Clearly, besides degree-first evaluation, we also have other evaluation strategies, such as depth-first.
+An interesting strategy is optimistic region evaluation, where we evaluate all region spaces in parallel, but before doing the slice transformations on a region, we check whether the region can be skipped using the Apriori property.
+Selecting the optimal strategy for performance is crucial in implementing multi-relational algebra, but a detailed exploration is beyond this paper’s scope and will be addressed in future work.
+4.3 Main application categories
 We briefly summarize three representative categories of applications:
@@
     Experimentation analysis. Several experimentation platforms have been built to slice the data to do A/B testing to understand the fine-grained impact of an experiment.
-The multi-relational algebra pipes service provides high-level abstractions that benefit all these applications:
-Our approach drastically simplifies system construction.
-For instance, in experimentation analysis platforms, teams often develop custom slicing-analysis kernels, each with tens of thousands of lines of code, typically built on Beam for its flexibility over SQL.
-Our multi-relational algebra pipes streamline these kernels into a concise pipeline of operators, enhancing interoperability with SQL due to their shared relational data foundation.
-This results in significant simplification of both experimentation analysis and integration with relational databases.
-Enabling Advanced Capabilities.
-The composability of multi-relational operators has unlocked new and powerful functionalities.
-For example, in system monitoring, techniques like “recurrent crawl,” 3.5 built on these operators, have achieved superior precision and recall in anomaly detection compared to traditional statistical methods.
-This demonstrates the ability of multi-relational operators to express complex analysis patterns that would be difficult to achieve otherwise.
-Enhanced Capabilities, Reduced Costs.
-The flexibility of slice transformations enables seamless integration of the Python Data Science stack with multi-relational processing.
-This allows for advanced analysis, such as quickly identifying revenue anomalies and their correlation with campaign budget allocation, leading to significant revenue gains in some cases.
-Additionally, migrating ad-hoc experimentation analysis systems to multi-relational pipes not only expands their capabilities to analyze finer-grained regions but also surprisingly results in improved latency, scalability, and reduced resource costs.
-This efficiency stems from the unified logical design of multi-relational algebra, enabling systematic optimizations across diverse analytical tasks.
+The multi-relational pipe service provides useful abstractions that benefit all these applications:
+Simplified systems.
+In experimentation analysis platforms, teams often build custom slicing-analysis kernels, each involving tens of thousands of lines of code, typically using Apache Beam [Pyt] due to its flexibility over SQL.
+Our multi-relational algebra pipes streamline these into concise multi-relational pipes.
+These pipelines, sharing a relational data foundation with SQL, improve interoperability.
+The outcome is substantial simplification for both experimentation analysis and integration with relational databases.
+Enhanced Capabilities.
+The ability to compose multi-relational operators has opened doors to powerful new functionalities.
+For instance, techniques such as “recurrent crawl” (Section 3.5) have leveraged these operators in system monitoring to achieve significantly better anomaly detection compared to traditional statistical methods alone.
+This showcases the power of multi-relational operators in expressing complex analysis patterns.
+Additionally, the flexibility of slice transformations allows for seamless integration with the Python Data Science stack, facilitating advanced analyses.
+In practice, this has enabled rapid identification of revenue anomalies and their correlation with campaign budgets, leading to substantial revenue increases in some cases.
+Reduced Costs.
+Migrating ad-hoc experimentation analysis systems to multi-relational pipes offers a twofold advantage: it empowers these systems to analyze data at a more granular level and, counter-intuitively, leads to improved latency, scalability, and reduced resource costs over the ad-hoc constructions.
+This efficiency gain arises from the unified logical design of multi-relational algebra, which allows for systematic optimizations across a wide range of analytical tasks.
 5 Conclusion
@@
 Note that we recognize the relation ‘employee’ as a multi-relation with a single schema.
-Appendix B Extended operators
+Appendix B Encoding a sparse relation
+To avoid the waste of encoding 0 in Section 3.1, we observe that schemas are the only important thing in a multi-relation to tell whether we have multiple relations, or just a single relation.
+To explain this, recall that a transaction set can certainly be encoded succinctly as a set of tuples with different schemas, namely like
+(Diaper=1,),
+(Diaper=1, Beer=1,),
+(Rum=1, Whiskey=1,),
+Now, instead of directly encoding this as a multi-relation with different schemas (thus this multi-relation indeed contains multiple relations of different schemas), we can encode this as a single relation, by just declaring a single schema that is the list of all items.
+Note that, the tuples can still have different schemas, but encoding that there is a single schema makes this into a relation with one schema, instead of a multi-relation.
+We can visualize this relation as follows, note that there is no 0 any more, but values are just “missing” (we don’t even need to pad NULL).
+Beer Diaper Rum Whisky Count
+$\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{1}$
+$\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{1}$
+$\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{1}$
+Table 3: Encoding a sparse table. We only use a wide schema to encode that this is a single relation.
+Now, when we perform MultiSelect, the schema matching process will only find one relation as matching, since there is only one schema!
+Further, for a particular region space, the semantics is simply tuples that have values on that region space.
+For example, for region space [Diaper], the slice representation will simply identifies tuples that have value $1$ on Diaper (since $1$ is the only value in this relation), and therefore there is only a single region (Device=1), which always corresponds to a legit pattern.
+Clearly, ‘support‘ can be easily computed from here.
+Appendix C Extended operators
 The operators described in the last section are fundamental, and already enabled a lot of otherwise-cumbersome analysis if one stays with RA.
 However, similar to that RA is extended with GroupBy, we consider two group-by extensions.
-B.1 MultiGroupBy
+C.1 MultiGroupBy
 In the RA, GroupBy partitions a relation into groups, and then uses an aggregation function to condense each group into a tuple, and finally we coalesce these tuples back into one relation.
-For BiRA, the situation is similar, we partition a bilevel relation into groups, where each group is a bilevel relation.
-Then we transform each group into another bilevel relation.
-Finally, we coalesce these bilevel relations into one by flattening the group key into the regions of each bilevel relation, and take union of all region-feature pairs.
+For Multi-Relational Algebra, the situation is similar, we partition a multi-relation into groups, where each group is a multi-relation.
+Then we transform each group into another multi-relation.
+Finally, we coalesce these multi-relations into one by flattening the group key into the regions of each multi-relation, and take union of all multi-relations.
 Definition 16 (MultiGroupBy).
-BiGroupBy is an operator as follows:
-BiGroupBy(
-br,
+MultiGroupBy is an operator as follows:
+MultiGroupBy(
+$\Psi$,
 by_attributes,
-bilevel_transform,
+multi_relational_transform,
 This works as follows:
   1. 1.
-    We partition br using by_attributes.
+    We partition $\Psi$ using by_attributes.
   2. 2.
-    For each group, we apply bilevel_transform to transform each bilevel relation to another.
+    For each group, we apply multi_relational_transform to transform each multi-relation to another.
   3. 3.
-    We flatten the group keys into the regions of t bilevel relation, and take union of all region-feature pairs.
-B.2 CoMultiGroupBy
-Consider two bilevel relations br1 and br2, and we groupby for each on Country.
-Now, we want the bilevel relation in br1 under Country=US to co-occur with the bilevel relation in br2 under Country=US.
+    We flatten the group keys into the corresponding multi-relations, and take a union of all multi-relations.
+C.2 CoMultiGroupBy
+Consider two multi-relations $\Psi_{1}$ and $\Psi_{2}$, and we groupby for each on Country.
+Now, we want the multi-relation in $\Psi_{1}$ under Country=US to co-occur with the multi-relation in $\Psi_{2}$ under Country=US.
 Such a need of co-occurrence is not new: this also appears in Apache Beam CoGroupByKey [CoG].
-However, unlike Beam which only needs to work with a data model of general dictionaries, here we must eventually output one bilevel relation.
-Suppose we use br1{Country=US} to denote the group Country=US after groupby, then co-groupby will put this together with br2{Country=US}, so we have a pair of bilevel relations (br1{Country=US}, br2{Country=US}).
-Then, we can use a BiJoin to condense this pair of bilevel relations back into one.
-After this, we are then in the situation of BiGroupBy, which we know how to handle.
+However, unlike Beam which only needs to work with a data model of general dictionaries, here we must eventually output one multi-relation.
+Suppose we use $\Psi_{1}${Country=US} to denote the group Country=US after groupby, then co-groupby will put this together with $\Psi_{2}${Country=US}, so we have a pair of multi-relations ($\Psi_{1}${Country=US}, $\Psi_{2}${Country=US}).
+Then, we can use a MultiJoin to condense this pair of multi-relations back into one.
+After this, we are then in the situation of MultiGroupBy, which we know how to handle.
 Definition 17 (CoMultiGroupBy).
-CoBiGroupBy is an operator as follows:
-CoBiGroupBy(
-br1,
-br2,
+CoMultiGroupBy is an operator as follows:
+CoMultiGroupBy(
+$\Psi_{1}$,
+$\Psi_{2}$,
 by_attributes,
-bijoin,
+multi_join,
 This works as follows:
   1. 1.
-    We partition br1 and br2 by by_attributes to form groups, where each group is a bilevel relation.
+    We partition $\Psi_{1}$ and $\Psi_{2}$ by by_attributes to form groups, where each group is a multi-relation.
   2. 2.
-    For each group key K, we have a bilevel relation br1{K}, and another bilevel relation br2{K}.
+    For each group key $K$, we have a multi-relation $\Psi_{1}\{K\}$, and another multi-relation $\Psi_{2}\{K\}$.
   3. 3.
-    We join these two bilevel relations into one, via bijoin(br1{K}, br2{K}).
-After the above steps, we are in the same situation as BiGroupBy, which we handle accordingly.
+    We join these two multi-relations into one using multi_join($\Psi_{1}\{K\}$, $\Psi_{2}\{K\}$).
+After the above steps, we are in the same situation as MultiGroupBy, which we handle accordingly.

### v4

@@
 That is, operators in the multi-relational algebra are about processing slices encoded in a multi-relation.
+For this, a multi-relational operator will first represent a multi-relation as a set of slices.
 We start with the definition of slice.
@@
 A multi-relational block, or simply block, is a pair of schemas $(\Gamma,\operatorname{{\mathcal{}F}})$, where $\Gamma\cap\operatorname{{\mathcal{}F}}=\emptyset$.
-Definition 4 (Representing a block in a multi-relation).
+We now define a first slice representation method using minimal schema matching. This representation method is mostly commonly used in our current applications.
+Definition 4 (Slice representation of a block via Minimal Schema Matching).
 Let $\Psi$ be a multi-relation, and $(\Gamma,\operatorname{{\mathcal{}F}})$ be a block.
@@
 Finally, we partition $Z|_{\operatorname{{\mathcal{}T}}}$ over $\Gamma$, and get a set of slices $(r,X)$, where $schema(r)=\Gamma$ and $schema(X)=\operatorname{{\mathcal{}F}}$.
-For (1), if there are multiple relations satisfy that condition, there are multiple ways to handle tie (for example, we can break the tie arbitrarily, or simply report an error).
-We leave details of handling ties later.
+A visualization of a slice representation is shown in Table 2.
+Region $\operatorname{{\mathcal{}F}}$
+$r_{1}$ $X_{1}$
+$r_{2}$ $X_{2}$
+Table 2: Slice representation of a block for a multi-relation.
+We have a region space $\Gamma$ and one feature set $\operatorname{{\mathcal{}F}}$.
+The slice representation is organized as a table where each row is a slice.
+Note that for Minimal Schema Matching, it might happen that for a general multi-relation, there are multiple relations satisfy that condition.
+In our applications, this is usually not a concern since they mostly want to work on the results of a previous groupby grouping sets query.
+Note also that there are multiple ways to handle tie (for example, we can break the tie arbitrarily, or simply report an error).
+There are cases where we need more stringent ways than Minimal Schema Matching for slice representation, which we describe in Appendix A.1.
 An important use of representation is to represent a multi-relation as different sets of slices. Below is an example
@@
 feature_set=[Revenue].
+With block representation, we now define slice representation for multiple region spaces and one feature set. In Appendix A.2, we will generalize this to multiple feature sets.
+Definition 5 (Slice representation of a multi-relation with multiple region spaces and one feature set).
+Let $\Gamma_{1},\dots,\Gamma_{N}$ be region spaces, and $\operatorname{{\mathcal{}F}}$ be one feature set.
+A slice representation of a multi-relation $\Psi$ w.r.t.
+the region spaces and feature sets is a table where each row is a slice.
+Specifically, row $i$ is of the form $(r_{i},X_{i})$, where $r_{i}$ is a region tuple of schema of some region space, and $X_{i}$ is a feature table of schema $\operatorname{{\mathcal{}F}}$
 Before we proceed to discuss operators, we define a semi-lattice structure on tuples which turns out to be important when processing regions.
@@
 We will later generalize and consider 2-slice transformation (i.e., which takes two slices as input), which are most useful for our applications.
-Definition 7 (MultiSelect).
-Let $\Psi$ be a multi-relation, $\Gamma_{1},\dots,\Gamma_{n}$ be $n$ region space schemas, $M_{1},\dots,M_{k}$ be $k$ slice transformations, where $M_{i}$ has feature set $\operatorname{{\mathcal{}F}}_{i}$, and results_predicates be predicates for filtering results.
-A MultiSelect is defined as
+We are ready to define MultiSelect.
+For the main text we will define MultiSelect for one slice transformation.
+Appendix B generalizes this to multiple slice transformations with different feature sets.
+Definition 8 (MultiSelect with one slice transformation).
+Let $\Psi$ be a multi-relation, $\Gamma_{1},\dots,\Gamma_{n}$ be $n$ region space schemas, $M$ be a slice transformation which requests feature set $\operatorname{{\mathcal{}F}}$.
+Finally, selection_condition is a predicate for filtering results.
+A MultiSelect is defined as:
 MultiSelect(
@@
 region_spaces=[$\Gamma_{1}$, ..., $\Gamma_{n}$]
-slice_transformations=[$M_{1}$, ..., $M_{k}$],
-results_predicates,
-and is processed as follows: For each $i\in[n],j\in[k]$,
+slice_transformations=[$M$],
+selection_condition,
+The evaluation of MultiSelect is as follows:
   1. 1.
-    For each $\Gamma_{i}$ and $\operatorname{{\mathcal{}F}}_{j}$ (of model $M_{j}$) we compute the slice representation $\operatorname{{\mathcal{}S}}^{i}_{j}$ of $\Psi$ w.r.t.
-$(\Gamma_{i},\operatorname{{\mathcal{}F}}_{j})$.
+    Slice representation.
+Given the region spaces $\Gamma_{1},\dots,\Gamma_{n}$ and feature sets $\operatorname{{\mathcal{}F}}$, we compute a slice representation $\rho$ of $\Psi$, where row $i$ of the slice representation $\rho$ is a slice $(r_{i},X_{i})$, where $X_{i}$ has schema $\operatorname{{\mathcal{}F}}$.
   2. 2.
-    We apply $M_{j}$ to all slices in $\operatorname{{\mathcal{}S}}^{i}_{j}$, and filter the output slices using results_predicates.
+    Slice transformation and output slice representation. For row-$i$, we evaluate each slice transformation $M$ and compute an output slice $(r^{\prime}_{i},X^{\prime}_{i})$.
+  3. 3.
+    Filter slice rows. The selection_condition is used to filter each output slice. (specifically, if $X_{i}^{\prime}$ is filtered as an empty relation, the slice is filtered).
+  4. 4.
+    Output multi-relation.
+We group output slices according to their schemas (the only schema non-uniformity comes from region spaces from our assumptions), and each group of slices are union into one relation.
+The result is a multi-relation, which serves as the output.
 Example 2.
-Let $\Psi$ be a multi-relation as specified in Figure 2, LocalTimeSeriesOutlierDetect be 1-slice transformation which detects timeseries outlier with feature set [Date, SumRevenue] (which computes a p_val for every date in a date range), results_predicates be ‘p_val < .001’, and region spaces be [Device] and [Browser].
+Let $\Psi$ be a multi-relation as specified in Figure 2, LocalTimeSeriesOutlierDetect be 1-slice transformation which detects timeseries outlier with feature set [Date, SumRevenue].
+Specifically, given a timeseries, it detects whether there is an outlier for a given date, and outputs a p_val for that date to indicate the strength of evidence that it is an outlier.
+Finally, let selection_condition be ‘p_val < .001’, and region spaces be [Device] and [Browser].
 The following MultiSelect performs outlier detection for [Date, SumRevenue] timeseries slices for each specified region space:
@@
 time_attr=Date,
-metric_attr=SumRevenue
-predicates=[p_val < .001,]
+metric_attr=SumRevenue,
+detection_date=’2024-09-01’,
+selection_condition=[p_val < .001],
 Specifically:
-  • for region space [Device], we match the relation with schema [Device, Date, SumRevenue] in $\Psi$, then we partition by Device, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and finally filter the output slice using ‘p_val < .001‘.
-  • for region space [Browser], we match the relation with schema [Browser, Date, SumRevenue] in $\Psi$, then we partition by Browser, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and finally filter the output slice using ‘p_val < .001‘.
+  • for region space [Device], we match the relation with schema [Device, Date, SumRevenue] in $\Psi$, then we partition by Device, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and detect outliers for the date 2024-09-01.
+We filter the output with ‘p_val < .001‘, thus only finding devices for which the p-value on 2024-09-01 is less than .001.
+  • for region space [Browser], we match the relation with schema [Browser, Date, SumRevenue] in $\Psi$, then we partition by Browser, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and detect outliers for the date 2024-09-01.
+We filter the output with ‘p_val < .001‘, thus only finding browsers for which the p-value on 2024-09-01 is less than .001.
 MultiSelect generalizes GROUP-BY GROUPING SETS and WINDOW functions.
 If we recognize a usual relation as a multi-relation with a single schema, then we observe that MultiSelect generalizes GROUP-BY GROUPING SETS.
-This is because GROUP-BY GROUPING SETS is about slice transformations that aggregate a single column.
-Therefore, MultiSelect also trivially generalize GROUP-BY and CUBE [GBLP96], which are special cases of GROUP-BY GROUPING SETS.
-MultiSelect also generalizes WINDOW functions, since we can use a slice transformation working on both the window column and metric column, and computes window-based aggregation (a more detailed explanation for this is deferred to A).
+To see this, note that each grouping set can be treated as a region space, and aggregation of several columns can be captured by a slice transformation that requests those columns as a feature set, and aggregate each column using desired aggregation functions.
+For example, suppose we want to do aggregations SUM(Revenue) and MAX(Clicks), then we can request a feature set [Revenue, Clicks], and the slice transformation performs sum and max aggregation on the respective column.
+With this, MultiSelect also generalizes CUBE [GBLP96], which are special cases of GROUP-BY GROUPING SETS.
+Note that in these cases, the slice transformation outputs a single tuple as output.
+MultiSelect also generalizes WINDOW functions, where we can use a slice transformation working on both the window column and metric column, and computes window-based aggregation, which in this case the output is a table, not a tuple.
+The details are deferred to Appendix C.
 Finally, we note that MultiSelect is trivially more general than all these operators since it can work on multi-relations.
@@
 In the multi-relational algebra, MultiProject restricts a multi-relation to a set of schemas.
-Formally, MultiProject applies the block representation procedure 4 to multiple blocks:
+Formally, MultiProject applies the block representation procedure 4 to multiple blocks induced by different region spaces (similar to MultiSelect, for the main text we only consider a single feature set; and the case for multiple feature sets can be handled using similar ideas described in Appendix B)
 Definition 12 (MultiProject).
-Let $\Gamma_{1},\dots,\Gamma_{n}$ be region space schemas, and $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{k}$ be feature sets, then the MultiProject operator
+Let $\Gamma_{1},\dots,\Gamma_{n}$ be region space schemas, and $\operatorname{{\mathcal{}F}}$ be feature sets, then the MultiProject operator
 MultiProject(
@@
 region_spaces=[$\Gamma_{1},\Gamma_{2},\dots,\Gamma_{n}$],
-feature_sets=[$\operatorname{{\mathcal{}F}}_{1},\operatorname{{\mathcal{}F}}2,\dots,\operatorname{{\mathcal{}F}}_{m}$],
-is the union of blocks for all $(\Gamma_{i},\operatorname{{\mathcal{}F}}_{j})$.
+feature_sets=[$\operatorname{{\mathcal{}F}}$],
+computes a slice representations for all blocks $\{(\Gamma_{i},\operatorname{{\mathcal{}F}})\ |\ i\in[n]\}$, and then convert the slice representation to an output multi-relation, same as described in MultiSelect (Definition 8).
 With MultiSelect and MultiProject we can define CRAWL operator:
@@
 and the merge is captured by a 2-slice transformation.
+We formalize this below.
+Similar to MultiSelect and MultiProject, we only define the case for one slice transformation, which can be generalized to multiple slice transformations using ideas described in Appendix B.
 Definition 14 (MultiJoin).
@@
 region_join_conditions,
-slice_transformations=[$M_{1},\dots,M_{k}$],
-Suppose we associate a region $r$ in $\Psi_{1}$ with another region $r^{\prime}$ in $\Psi_{2}$ that are joinable w.r.t.
-the specified conditions, For any 2-slice transformation $M_{i}$, it has two features sets $\operatorname{{\mathcal{}F}}_{1}$, $\operatorname{{\mathcal{}F}}_{2}$, from which we can identify a slice $S=(r,X)$ and $S^{\prime}=(r^{\prime},X^{\prime})$, and so we can compute a slice using $M(S,S^{\prime})$.
-We repeat this for all models, and merge all output slices into a multi-relation.
+slice_transformations=[$M$],
+where the 2-slice transformation $M$ has two features sets $\operatorname{{\mathcal{}F}}_{1}$, $\operatorname{{\mathcal{}F}}_{2}$, and $\Gamma_{1},\dots,\Gamma_{n}$ are region spaces.
+For evaluating this operator, we do a slice representation $\rho_{1}$ for $\{(\Gamma_{i},\operatorname{{\mathcal{}F}}_{1})\ |\ i\in[n]\}$, and another slice representation $\rho_{2}$ for $\{(\Gamma_{i},\operatorname{{\mathcal{}F}}_{2})\ |\ i\in[n]\}$.
+Then for a slice $\theta=(r,X)\in\rho_{1}$ and $\theta^{\prime}=(r^{\prime},X^{\prime})\in\rho_{2}$, if regions $r$ and $r^{\prime}$ are joinable w.r.t.
+the specified conditions, then we compute an output slice using $M(\theta,\theta^{\prime})$.
+The final result is an output slice representation, which we convert to a multi-relation same as in MultiSelect (Definition 8).
 One application of MultiJoin is the Inner-Outer Join and recurrent crawl, which we describe in Section 3.4.
@@
     Astute readers may argue that our approach requires encoding transactions in a very wide table, with most of the table 00 (the number of columns is the number of distinct items), and it is wasteful to do this encoding.
-There is an elegant way to avoid such waste by exploiting the schema of a multi-relation, which we discuss in Section B.
+There is an elegant way to avoid such waste by exploiting the schema of a multi-relation, which we discuss in Section D.
 Id Transaction
@@
 Mutl-Relational Algebra significantly broadens the scope of algebraic theory for relational analytics, and has found various applications in data insights.
-Appendix A Emulating WINDOW functions
+Appendix A More on data modeling
+A.1 More on slice representations
+We describe two methods for more rigid slice representations. Both methods require more information from the user, passing as parameters of a multi-relational operator.
+Rigid Schema Matching.
+In the Minimal Schema Matching method, we ask the user to pass in a list of feature sets $[\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{k}]$.
+In the Rigid Schema Matching, for each feature schema we ask the user to pass in an additional “source feature schema”, so it is like $[(\operatorname{{\mathcal{}F}}_{1},\widehat{\operatorname{{\mathcal{}F}}}_{1}),\dots,(\operatorname{{\mathcal{}F}}_{k},\widehat{\operatorname{{\mathcal{}F}}}_{k})]$.
+At the time of slice representation, for each region space $\Gamma_{i}$ and $(\operatorname{{\mathcal{}F}}_{j},\widehat{\operatorname{{\mathcal{}F}}}_{j})$, we will construct schema $\widehat{S}_{ij}=\Gamma_{i}\cup\widehat{\operatorname{{\mathcal{}F}}}_{j}$, and find a relation $R$ in the multi-relation with schema exactly $\widehat{S}_{ij}$.
+If there are multiple matching, it is an error.
+From $R$, then we project to $\operatorname{{\mathcal{}F}}_{j}$ for features.
+For readers familiar with SQL, one can think of $\widehat{F}_{j}$ as providing a FROM clause to indicate the relation we want to work on.
+Minimal Containing Dimensions Schema Matching.
+In the Minimal Containing Dimensions Schema Matching, we ask the user to additionally pass in a list $D$ of dimension attributes (which is a subset of all attributes).
+Now given a region schema $\Gamma_{i}$ and feature set $\operatorname{{\mathcal{}F}}_{j}$ (let $S_{ij}=\Gamma_{i}\cup\operatorname{{\mathcal{}F}}_{j}$, the desired relation is identified as follows: We first go through all relations that contain the schema $S_{ij}$, and pick relations such that their dimension set (intersecting with $D$) is minimal.
+Then among all these relations, we pick the one with the minimal schema.
+For ties, we can either break the tie arbitrarily, or report an error.
+A.2 Generalizing slice representation to multiple feature sets
+Definition 17 (Slice representation of a multi-relation with multiple region spaces and multiple feature sets).
+Let $\Gamma_{1},\dots,\Gamma_{N}$ be region spaces, and $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{M}$ be feature sets.
+A slice representation of a multi-relation $\Psi$ w.r.t.
+the region spaces and feature sets is a table where each row is called a slice tuple or slice row.
+Specifically, row $i$ is of the form $(r_{i},X^{i}_{1},X^{i}_{2},\dots,X^{i}_{M})$, where $r_{i}$ is a region tuple of schema of some region space, and $X^{i}_{j}$ is a feature table of schema $\operatorname{{\mathcal{}F}}_{j}$ ($j\in[M]$).
+A visualization of a slice representation is shown in Table 4.
+Region $\operatorname{{\mathcal{}F}}_{1}$ $\operatorname{{\mathcal{}F}}_{2}$cdots $\operatorname{{\mathcal{}F}}_{M}$
+$r_{1}$ $X^{1}_{1}$ $X^{1}_{2}$cdots $X^{1}_{M}$
+$r_{2}$ $X^{2}_{1}$ $X^{2}_{2}$cdots $X^{2}_{M}$
+Table 4: Slice representation of a multi-relation.
+We have multiple region spaces $\Gamma_{1},\dots,\Gamma_{N}$, and multiple feature sets $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{M}$.
+The slice representation is organized as a table of multiple slice tuples (or slice rows).
+Each row consists of a region tuple, followed by $M$ feature tables, one for each schema $\operatorname{{\mathcal{}F}}_{j}$, $j\in[M]$.
+Appendix B Generalizing MultiSelect to multiple feature sets
+B.1 Multiple tuple-output slice transformations
+We now define MultiSelect with multiple tuple-output slice transformations (instead of one table-ouput slice transformation).
+Definition 18 (MultiSelect with multiple tuple-output slice transformations).
+Let $\Psi$ be a multi-relation, $\Gamma_{1},\dots,\Gamma_{n}$ be $n$ region space schemas, $M_{1},\dots,M_{k}$ be $k$ slice transformations, where $M_{i}$ requests feature set $\operatorname{{\mathcal{}F}}_{i}$ and is region-invariant.
+Further, the output feature schemas of slice transformations are pair-wise disjoint.
+For simplicity, we assume that on input $(r,X)$, $M_{i}$ outputs $(r,\zeta)$, where $\zeta$ is a single tuple.
+Finally, selection_condition is a predicate for filtering results (the specification is the same as selection condition of SELECT operator in the relational algebra; we explain this below).
+A MultiSelect is defined as:
+MultiSelect(
+$\Psi$,
+region_spaces=[$\Gamma_{1}$, ..., $\Gamma_{n}$]
+slice_transformations=[$M_{1}$, ..., $M_{k}$],
+selection_condition,
+The evaluation of MultiSelect is as follows:
+  1. 1.
+    Slice representation.
+Given the region spaces $\Gamma_{1},\dots,\Gamma_{n}$ and feature sets $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{k}$, we compute a slice representation $\rho$ of $\Psi$, where row $i$ of the slice representation $\rho$ is a slice tuple $(r_{i},X^{i}_{1},\dots,X^{i}_{k})$, where $X^{i}_{j}$ has schema $\operatorname{{\mathcal{}F}}_{j}$.
+  2. 2.
+    Slice transformation and output slice representation.
+For a slice row, we evaluate each slice transformation $M_{j}$ on each of the slice $(r_{i},X^{i}_{j})$, and compute output slice $(r_{i},\zeta^{i}_{j})$.
+The result is another slice representation, where row-$i$ is $(r_{i},\zeta^{i}_{1},\dots,\zeta^{i}_{k})$.
+  3. 3.
+    Filter slice rows.
+The selection_condition is used to filter each slice tuple of the output from the previous step.
+Specifically, since $r_{i}$ and $\zeta^{i}_{j}$ are all tuples, all tuples can be concatenated into one tuple, so selection_condition applies to filter each tuple.
+(this is why selection_condition can be specified the same way as in the relational algebra)
+  4. 4.
+    Output multi-relation.
+We group tuples according to their schemas (the only schema non-uniformity comes from region spaces from our assumptions), and each group of tuples are union into one relation.
+The result is a multi-relation, which serves as the output.
+B.2 Multiple table-output slice transformations
+[…]
+In this section, we investigate the question of slice representation output where there are multiple slice transformations, and each output a feature table.
+In particular, we relax the requirement and allow output schemas of different slice transformations to have overlap.
+For simplicity, we still require that all slice transformations are region-invariant, though the argument can be carried over to allow slice transformation that are not region-invariant.
+We now consider the general situation.
+We have an output slice row of the form $(r,Z_{1},Z_{2},\dots,Z_{k})$, where $Z_{i}$ is the output feature table of the slice transformation $M_{i}$.
+Further, $schema(Z_{i})\cap schema(Z_{j})$ may be non-empty.
+The first question is why do we want to merge these feature tables?
+Can’t we just have a multi-relation where we create one schema for each $(r,Z_{i})$?
+The first motivation is exactly as in the case join, we want to meld fragmented information together, as much as possible.
+Second, following the first motivation, a user may want to access two output columns, one from $Z_{i}$ and one from $Z_{j}$, and we need that there is one single relation in the multi-relation to hold both columns.
+One solution to this question is to introduce a new concept, called dimension, into the theory.
+A dimension is an attribute which serves the role of determining whether two tables can be merged via join.
+Only a subset of attributes may be dimensions.
+Specifically, we would join two feature tables $Z_{i}$ and $Z_{j}$ if and only if they share the same dimension set.
+When that happens, we do a full outer-join on the dimension columns.
+Specifically, we say that $Z_{i}\sim Z_{j}$, if $dims(Z_{i})=dims(Z_{j})$.
+$\sim$ is an equivalence relation, and so we can partition $Z_{1},\dots,Z_{k}$ into equivalence classes, and within each we can join.
+With this, therefore we can output an output slice row $(r,Z^{\prime}_{1},\dots,Z^{\prime}_{q})$, where $q$ is the number of equivalence classes.
+Therefore, dimensions allow us to merge information as much as possible, as indicated by the user.
+To implement this, we extend the MultiSelect operator with an additional output_feature_dimensions parameter, which allows the user to configure what are the all possible output feature dimensions.
+We give a concrete example of this.
+Suppose that we have three slice transformations.
+The first performs a smoothing on features [Date, Cost].
+The second performs a different smoothing on features [Date, Clicks].
+The third computes a CostPerClick on features [Cost, Clicks].
+Suppose that the input is a multi-relation $\Psi$ with two relations, the first relation $R_{1}$ with schema [Date, Cost, Clicks], and the second $R_{2}$ with schema [Cost, Clicks].
+This multi-relation may be obtained using a group-by grouping sets query, such as SELECT Date, SUM(Cost), Sum(Clicks) FROM r GROUP BY GROUPING SETS [], [Date].
+Note that the region space is empty here.
+Therefore, a slice row looks like $(r=(),Z_{1},Z_{2},Z_{3})$ where $Z_{1}$ is a table on [Date, Cost], $Z_{2}$ is a table on [Date, Clicks], and $Z_{3}$ is a tuple (CostPerClick=XXX).
+A user may naturally want to join $Z_{1}$ and $Z_{2}$, since later we want to analyze Cost and Clicks together.
+With dimensions, the user can then configure output_feature_dimensions to be [Date].
+Then we have two equivalence classes, $[Z_{1},Z_{2}]$ and $[Z_{3}]$, and we will join the tables in the first equivalence class on Date and get $Z_{1}^{\prime}$, which has schema [Date, Cost, Clicks].
+The output final slice row is $(r=(),Z_{1}^{\prime},Z_{2}^{\prime}=Z_{3})$.
+Appendix C Emulating WINDOW functions
 A typical WINDOW function query is as follows:
@@
 Then for each partition, we sort by Age, and compute a moving average of Salary for each row.
-We can emulate this using MultiSelect as follows:
+We can emulate this using MultiSelect (with a single slice transformation but which outputs a table) as follows:
 MultiSelect(
@@
 region_spaces=[[Department]],
-slice_functions=[
+slice_transformations=[
 MOVING_AVERAGE(
@@
 Note that we recognize the relation ‘employee’ as a multi-relation with a single schema.
-Appendix B Encoding a sparse relation
+Appendix D Encoding a sparse relation
 To avoid the waste of encoding 0 in Section 3.1, we observe that schemas are the only important thing in a multi-relation to tell whether we have multiple relations, or just a single relation.
@@
 Clearly, ‘support‘ can be easily computed from here.
-Appendix C Extended operators
+Appendix E Extended operators
 The operators described in the last section are fundamental, and already enabled a lot of otherwise-cumbersome analysis if one stays with RA.
 However, similar to that RA is extended with GroupBy, we consider two group-by extensions.
-C.1 MultiGroupBy
+E.1 MultiGroupBy
 In the RA, GroupBy partitions a relation into groups, and then uses an aggregation function to condense each group into a tuple, and finally we coalesce these tuples back into one relation.
@@
     We flatten the group keys into the corresponding multi-relations, and take a union of all multi-relations.
-C.2 CoMultiGroupBy
+E.2 CoMultiGroupBy
 Consider two multi-relations $\Psi_{1}$ and $\Psi_{2}$, and we groupby for each on Country.